# EPFL, Lausanne, Switzerland

### v1

@@ 2 Related Work
Instead of improving UDF-based schemes, one can also propose different representation schemes for open surfaces. In 3PSDF [4], they are modeled by using occupancy grids, by masking out empty regions, and by classifying unmasked grid cells as either positive or negative. These pseudo signs provide a notion of "inside" or "outside" the local surface based on the surface gradient of the nearest surface point, as in [21]. However, this method requires pre-orienting the training data, and does not provide the continuous distances required for precise meshing. Instead of directly learning the UDF, the CSP algorithm of [19] predicts the closest surface point. This representation is differentiable at the surface, but the scalar field does not exhibit distance-like continuity and is challenging to learn. In contrast, our proposed GDF predicts the vector pointing to the nearest surface point. When crossing the surface: like SDFs, GDFs scalar field changes from positive to negative or vice versa, while CSP yields constant values. As a result, CSPs are harder to train and yield less precise reconstructions, as we will demonstrate. One approach that is related to our method and CSP is NVF[23], which predicts the gradient vector for each query point. However, this method is specifically designed for 3D shape reconstruction from point-clouds where the input feature is constructed via neighboring points in the given point-cloud.
@@ 3.1 Definition
GDF : $\textrm{GDF : }\mathbb{R}^{3}$ $\rightarrow\mathbb{R}^{3},$
$\mathbf{x}\rightarrow\mathbf{v}$ $=\hat{\mathbf{x}}-\mathbf{x}\;$ , UNKNOWN (1)
@@
$u$ $=\|\mathbf{v}\|\;$ , UNKNOWN (2)
[…]
@@ 3.2 GDFs as Output of a Deep Network
$S$ =$=\{S_{i}=(\mathbf{x}^{i},\mathbf{v}^{i}),1\leq i\leq N\}\;,$
$\mathbf{x}^{i}$ =$=\{\mathbf{x}_{1}^{i}\ldots\mathbf{x}_{k}^{i}\}\;,$ (4)
$\mathbf{v}^{i}$ =$=\{\mathbf{v}_{1}^{i}\ldots\mathbf{v}_{k}^{i}\}\;,$
@@ 4.1 Datasets, Metrics, Baselines, and Settings
We compare our method against three other implicit-surface methods, NDF [7] that predicts UDF values directly, CSP [21] that regresses the closest surface point, and 3PSDF [4] that labels points as either “$-1$”,“$1$”, or “null”. To ensure fair comparison in each experiment, we train different networks with the same architecture, training data, and optimization settings. The only difference is how the output is representated. For each training instance, we extract 400000 points near the surface and 20000 points uniformly sampled throughout the entire volume. We use Adam optimizer [10] with a starting learning rate of $10^{-4}$ that decreases by 25% after each quater of the training. We use MeshUDF [9] as the meshing algorithm for all methods except for the 3PSDF representation that can be meshed using Marching Cubes. MeshUDF reasons on the gradient of the UDF for meshing. For NDF, we compute this gradient by running a back-propagating pass through the network, as in deepSDF [17]. For CSP and our method, the gradient can be computed directly from the output representation. In our case, it is given by Eq. 3, without the need to back-propagate. We never perform any post-processing step.
@@ 4.2 Representing a Single Complex Shape
In Table 1, we report the Chamfer Distances ×$\times 10^{4}$ (CD) and the Normal Consistency of the resulting surfaces from different methods. Our proposed GDF produces smooth meshes that closely match the ground-truth ones in all cases, as shown in Fig. 3. This is because surfaces in our GDF representation locate in continuous, differentiable parts of the scalar field, enabling the deep network to easily memorize them. On the contrary, the results of NDF exhibit staircase artifacts and surface holes, typical issues associated with the direct learning of UDFs. CSP creates smoother surfaces. However, they have many holes, despite being trained for an identical number of iterations as other methods. Compared to ours, this representation is more difficult to learn as it does not exhibit distance-like scalar field. The 3PSDF representation makes network training easier with fast convergence and no holes in the reconstructed surfaces, even for challenging cases such as the "reading room" of the 3D-Scene dataset shown in bottom row of the figure. However, this efficiency comes with a cost of discretizing distance values into either “-1” or “1,” resulting in less accurate meshes with noticeable staircase artifacts.
Table 1: Fitting to a single complex mesh. Average L2 Chamfer Distance ×$\times 10^{-4}$ (CD) and Normal Consistency (NC) for the MGN and 3D-Scenes datasets. We use the same MLP for each scene and only change the output representation.
@@ 4.3 Auto-Decoders to Model a Category of Shapes
Table 2: Representing known shapes using an auto-decoder. Average L2 Chamfer Distance ×$\times 10^{-4}$ (CD) and normal consistency (NC) for the MGN and ShapeNet-Car datasets.
@@
Table 3: Representing previously unseen shapes using an auto-decoder. Average L2 Chamfer Distance ×$\times 10^{-4}$ (CD) and normal consistency (NC) for the MGN and closed ShapeNet-Car datasets.
@@ 5.4 Gradient Distance Field Visualization
(a) Ground-Truth (b) Learned (c) Ground-Truth (d) Learned distance distance x-gradient x-gradient
@@
5.5 Limitations
While UDF or SDF only regress a distance value, GDF outputs a 3D vector at each query point. While we have shown that learning these three components in combination is still much easier than learning a UDF, learning an SDF is even simpler when dealing with closed surfaces. A risk with GDFs is the potential misalignment between sub-fields, that is, they may cross the zero-iso surface at different locations. In this case, the extracted distance might not reach zero. In practice, we have found this issue not to be significant. Small holes can occur, but only rarely. Another potential difficulty with learning a GDF is the non-differentiability at the medial axis of the object, which is the set of points having more than one closest point on the object surface. However, those points are often far away from the surface and inaccurate value estimations at these points usually do not affect the surface reconstruction accuracy.
@@ 6 Conclusion
Appendix 0.A Qualitative results on MGN dataset
(a) NDF[7] (b) CSP [19] (c) 3PSDF [4] (d) GDF (Ours) (e) GT
Figure 10: Representing known shapes from MGN[2] dataset using an auto-decoder. We use the same MLPs trained using different implicit representations.
(a) NDF[7] (b) CSP [19] (c) 3PSDF [4] (d) GDF (Ours) (e) GT
Figure 11: Representing known shapes from MGN[2] dataset using an auto-decoder. We use the same MLPs trained using different implicit representations.
Appendix 0.B Qualitative results on ShapeNet-Car dataset
(a) NDF[7] (b) CSP [19] (c) 3PSDF [4] (d) GDF (Ours) (e) GT
Figure 12: Representing known shapes from ShapeNet-Car(P) dataset using an auto-decoder. Car models are processed to be watertight and without inner-structures. We use the same MLPs trained using different implicit representations.
(a) NDF[7] (b) CSP [19] (c) 3PSDF [4] (d) GDF (Ours) (e) GT
Figure 13: Representing known shapes from ShapeNet-Car(R) dataset using an auto-decoder. We use the same MLPs trained using different implicit representations.
Appendix 0.C Comparison between MeshUDF[9] and DualMeshUDF[24]
We use MeshUDF[9] as the meshing algorithm for all methods except for 3PSDF that can be meshed using Marching-Cube[6, 11]. An alternative for meshing a UDF is a recent method of Zhang et al., DualMeshUDF[24]. Both MeshUDF and DualMeshUDF are very efficent, compared to the ball-pivoting algorithm[1]. We observe that DualMeshUDF are generally better at the boundary areas where MeshUDF often produces extra surfaces. However, the results of DualMeshUDF often exhibit many holes, as we show in Fig. 14, for all methods.
(a) NDF +[24] (b) NDF +[9] (c) GDF + [24] (d) GDF +[9] (e) GT
Figure 14: Car models reconstructed using MeshUDF[9] and DualMeshUDF[24]. DualMeshUDF yeilds meshes with accurate boundaries but with many holes.

### v2

# Gradient Distance Function

@@ 2 Related Work
Instead of improving UDF-based schemes, one can also propose different representation schemes for open surfaces. In 3PSDF [3], they are modeled by using occupancy grids, by masking out empty regions, and by classifying unmasked grid cells as either positive or negative. These pseudo signs provide a notion of "inside" or "outside" the local surface based on the surface gradient of the nearest surface point, as in [18]. However, this method requires pre-orienting the training data, and does not provide the continuous distances required for precise meshing. Instead of directly learning the UDF, the CSP algorithm of [16] predicts the closest surface point. This representation is differentiable at the surface, but the scalar field does not exhibit distance-like continuity and is challenging to learn. In contrast, our proposed GDF predicts the vector pointing to the nearest surface point. When crossing the surface: like SDFs, GDFs scalar field changes from positive to negative or vice versa, while CSP yields constant values. As a result, CSPs are harder to train and yield less precise reconstructions, as we will demonstrate. One approach that is related to our method and CSP is NVF [20], which predicts the gradient vector for each query point using codebooks to memorize UDF patterns that match the input point cloud. Our method, whereas, shows that directly learning gradient distance functions yields a general-purpose, surface-differentiable representation—well beyond codebook-style point-cloud modeling—while remaining simple to train with a standard auto-decoder.
@@ 3.1 Definition
$\textrm{GDF : }\mathbb{R}^{3}$ $\rightarrow\mathbb{R}^{3},$
$\mathbf{x}\rightarrow\mathbf{v}$ $=\hat{\mathbf{x}}-\mathbf{x}\;,$  (1)
@@
$u$ $=\|\mathbf{v}\|\;,$  (2)
$\mathbf{g}$ $=\frac{1}{u}\mathbf{v}\;.$  (3)
@@ 3.2 GDFs as Output of a Deep Network
$S$ $=\{S_{i}=(\mathbf{x}^{i},\mathbf{v}^{i}),1\leq i\leq N\}\;,$
$\mathbf{x}^{i}$ $=\{\mathbf{x}_{1}^{i}\ldots\mathbf{x}_{k}^{i}\}\;,$ (4)
$\mathbf{v}^{i}$ $=\{\mathbf{v}_{1}^{i}\ldots\mathbf{v}_{k}^{i}\}\;,$
@@ 4.1 Datasets, Metrics, Baselines, and Settings
We compare our method against three other implicit-surface methods, NDF [5] that predicts UDF values directly, CSP [18] that regresses the closest surface point, and 3PSDF [3] that labels points as either “1-1”,“11”, or “null”. To ensure fair comparison in each experiment, we train different networks with the same architecture, training data, and optimization settings. The only difference is how the output is representated. For each training instance, we extract 400000 points near the surface and 20000 points uniformly sampled throughout the entire volume. We use Adam optimizer [8] with a starting learning rate of $10^{-4}$ that decreases by 25% after each quater of the training. We use MeshUDF [7] as the meshing algorithm for all methods except for the 3PSDF representation that can be meshed using Marching Cubes. MeshUDF reasons on the gradient of the UDF for meshing. For NDF, we compute this gradient by running a back-propagating pass through the network, as in deepSDF [14]. For CSP and our method, the gradient can be computed directly from the output representation. In our case, it is given by Eq. 3, without the need to back-propagate. We never perform any post-processing step.
@@ 4.2 Representing a Single Complex Shape
In Table 1, we report the Chamfer Distances $\times 10^{4}$ (CD) and the Normal Consistency of the resulting surfaces from different methods. Our proposed GDF produces smooth meshes that closely match the ground-truth ones in all cases, as shown in Fig. 3. This is because surfaces in our GDF representation locate in continuous, differentiable parts of the scalar field, enabling the deep network to easily memorize them. On the contrary, the results of NDF exhibit staircase artifacts and surface holes, typical issues associated with the direct learning of UDFs. CSP creates smoother surfaces. However, they have many holes, despite being trained for an identical number of iterations as other methods. Compared to ours, this representation is more difficult to learn as it does not exhibit distance-like scalar field. The 3PSDF representation makes network training easier with fast convergence and no holes in the reconstructed surfaces, even for challenging cases such as the "reading room" of the 3D-Scene dataset shown in bottom row of the figure. However, this efficiency comes with a cost of discretizing distance values into either “-1” or “1,” resulting in less accurate meshes with noticeable staircase artifacts.
Table 1: Fitting to a single complex mesh. Average L2 Chamfer Distance $\times 10^{-4}$ (CD) and Normal Consistency (NC) for the MGN and 3D-Scenes datasets. We use the same MLP for each scene and only change the output representation.
@@ 4.3 Auto-Decoders to Model a Category of Shapes
Table 2: Representing known shapes using an auto-decoder. Average L2 Chamfer Distance $\times 10^{-4}$ (CD) and normal consistency (NC) for the MGN and ShapeNet-Car datasets.
@@
Table 3: Representing previously unseen shapes using an auto-decoder. Average L2 Chamfer Distance $\times 10^{-4}$ (CD) and normal consistency (NC) for the MGN and closed ShapeNet-Car datasets.
@@ 5.4 Gradient Distance Field Visualization
(a) GT distance (b) Learned distance Refer to caption (c) GT x-grad (d) Learned x-grad